OpenStudy (anonymous):

if i have room for 6 books on a bookshelf, how many different ways can i fit 9 different books on there?

OpenStudy (anonymous):

This is 9 permute 6.

OpenStudy (anonymous):

\[ _9P_6 = \frac{9!}{(9-6)!} = \frac{9!}{3!} = 9\cdot 8\cdot 7\cdot 6\cdot 5\cdot 4=60480 \]